OpenStudy (anonymous):

A fair ground ride spins its occupants inside a flying saucer-shaped container. if the horizontal circular path the riders follow has an 8.00 m radius, at how many revolutions per minute will the riders be subjected to a centripetal acceleration whose magnitude is 1.50 times that due to gravity?

OpenStudy (jacob902):

centripetal accel = 1.5*g ω²r = 1.5*9.8m/s² ω² * 8m = 14.7 m/s² ω = 1.36 rad/s * 1rev/2πrads * 60s/min = 12.9 rpm

OpenStudy (anonymous):

Thanks Jacob902. how did you get the 1.36rad/s